OpenStudy (anonymous):

If the area of a rectangle 24a2b and the length is 8ab2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer. the quantity of 3 times a all over b b over the quantity 3 times a 3 over the quantity a times b 2st

OpenStudy (anonymous):

\[\frac{ 3a }{ b }\] \[\frac{ b }{ 3a }\] \[\frac{ 3 }{ ab }\] 2st

OpenStudy (anonymous):

what problem are you facing here?

OpenStudy (anonymous):

If the area of a rectangle 24a^2b and the length is 8ab^2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer.

OpenStudy (anonymous):

and those are the answers

OpenStudy (anonymous):

i know the question, i want to know in which part you need help

OpenStudy (anonymous):

I have no idea how to do it in the first place lol

OpenStudy (anonymous):

simply divide 24a2b by 8ab2 as instructed! where are you facing problem?

OpenStudy (anonymous):

oh thts it?

OpenStudy (anonymous):

yeah!!! as simple as that :P

OpenStudy (anonymous):

I did not see tht... Thank You!

OpenStudy (anonymous):

welcome :P